(a) If the offered price of a Treasury bill with 74 days to maturity and a $10 million face amount is $9,916,306, what is the yield on a bank discount basis?

(b) If the bid yield on a bank discount basis is 1.5bp higher than the offered yield, what is the bid price?

(c) What is the bid-ask spread in dollars?
Answer #1

1.
=(1-9916306/10^6*1/10)*360/74=4.0716%

2.
=10*10^6*(1-(4.0716%+1.5%/100)*74/360)
=$9,915,997.67

3.
=9916306-9915997.67=308.33